# The Bayesian Origin of the Probability Weighting Function in Human Representation of Probabilities

### v1

@@ C.11.1 Application of BLO Model to DMR Task
Following Zhang et al. (2020), the estimate $\hat{\pi}(p)$ is integrated into Cumulative Prospect Theory (Tversky & Kahneman, 1992) to predict choice behavior; the certainty equivalent (CE) for a two-outcome lottery $(x_{1},p;x_{2},1-p)$ is given by:
@@
C.11.2 Reimplementation of BLO and LLO Models Variants
We first reimplemented the four main parametric models proposed by Zhang et al. (2020) for the DMR task. These included the BLO and LLO models, each combined with either a proportion-dependent variance compensation (V(p)) or a constant variance (Const V). As with the JRF task analysis, these models were fitted to the data for each subject individually. Our implementation utilized gradient-based optimizers (specifically, Adam or SignSGD) for parameter optimization.
@@ C.11.2 Reimplementation of BLO and LLO Models Variants
C.12 Cross-Task Analysis of Resources and Bias
C.12.1 Freely Fitted Prior and Resources in JRF and DMR tasks
The freely fitted priors and corresponding resources in three tasks are shown in Figure 33). Priors differ noticeably between tasks, but resources consistently exhibits the U-shape, with highest sensitivity near the extremes. Interestingly, in the DMR Choice tasks, the shape of the prior appears more closely aligned with the resources than in the other tasks. This prior–resources match is reminiscent of the account proposed by Frydman & Jin (2023), but the fact that it is not observed across all tasks suggests that such alignment is limited in generality.
@@ C.12.1 Freely Fitted Prior and Resources in JRF and DMR tasks
C.12.2 Bias Decomposition in JRF and DMR tasks
Across tasks, the decomposition of bias (Figure 34) shows that likelihood repulsion is consistent, whereas prior attraction varies more across conditions. This suggests that U-shaped encoding, which drives the repulsion effect, is a robust and shared property, while priors are more task-dependent.
Figure 34: Across Tasks: Bias decomposition (prior attraction, top; likelihood repulsion, bottom) across tasks. The solid line shows the group median, and the shaded area indicates the interquartile range (IQR). Plotting details are provided in Appendix C.13.
C.13 Details for Plotting Group-level Prior and Resources
@@
C.14 Methods for Calculating Bias and Variance
@@ C.14.3 Bayesian Model
C.15 Experiments on explaining Probability Distortion in Vision-Language Models
@@ C.15.1 Task and Stimuli
C.15.2 Behavioral Findings: S-shaped Distortion
We first examined the end-to-end behavior of open-source VLMs when prompted with dot-array images. Both InternVL3-14B and Qwen-VL2.5-7B produced probability estimates showing the classic S-shaped distortion observed in humans (Figures 35, 36). Specifically, models overestimated small probabilities, underestimated large probabilities, and showed a dip in accuracy near $p=0.5$, accompanied by relatively low variability—a pattern similar to human data. These results held across both long and short prompts (Figures 37, 38).
@@ C.15.2 Behavioral Findings: S-shaped Distortion
C.15.3 Representation Analysis with A Proxy for Fisher Information
@@
We applied this method to hidden states in the decoder part of each VLM, where vision and text features interact. Both InternVL3-14B and Qwen-VL2.5-7B exhibited robust U-shaped Fisher information profiles (Figures 40, 41). The U-shape was evident from the earliest fusion layers and remained stable through the final layers, with higher discriminability near $p=0$ and $p=1$. This pattern parallels our theoretical prediction that a U-shaped allocation of resources underlies S-shaped bias.
@@ B Layer 48
C.15.4 Summary
Together, these results show that S-shaped probability distortion also emerges in large VLMs when performing the JRF task. The distortion disappears under unimodal (text-only or simple image) inputs, but reappears under joint vision–language processing. Probing hidden representations reveals a U-shaped Fisher Information, consistent with the Bayesian account that nonuniform allocation of encoding resources underlies the observed S-shaped bias.

### v2

@@ C.12.1 Application of BLO Model to DMR Task
Following zhang2020, the estimate $\hat{\pi}(p)$ is integrated into Cumulative Prospect Theory (CITE) to predict choice behavior; the certainty equivalent (CE) for a two-outcome lottery $(x_{1},p;x_{2},1-p)$ is given by:
@@
C.12.2 Reimplementation of BLO and LLO Models Variants
We first reimplemented the four main parametric models proposed by zhang2020 for the DMR task. These included the BLO and LLO models, each combined with either a proportion-dependent variance compensation (V(p)) or a constant variance (Const V). As with the JRF task analysis, these models were fitted to the data for each subject individually. Our implementation utilized gradient-based optimizers (specifically, Adam or SignSGD) for parameter optimization.
@@ C.12.2 Reimplementation of BLO and LLO Models Variants
C.13 Cross-Task Analysis of Resources and Bias
C.13.1 Freely Fitted Prior and Resources in JRF and DMR tasks
The freely fitted priors and corresponding resources functions in three tasks are shown in Figure 32). Priors differ noticeably between tasks, but resources consistently exhibits the U-shape, with highest sensitivity near the extremes. Interestingly, in the DMR Choice tasks, the shape of the prior appears more closely aligned with the resources than in the other tasks. This prior–resources match is reminiscent of the account proposed by frydman2023source, but the fact that it is not observed across all tasks suggests that such alignment is limited in generality.
@@ C.13.1 Freely Fitted Prior and Resources in JRF and DMR tasks
C.13.2 Bias Decomposition in all tasks
As shown in Section 3.1, we decomposed the total estimation bias ($\mathbb{E}[\hat{p}|p]-p$) into three components: Prior Attraction, Likelihood Repulsion, and Boundary Regression. We computed these components by:
Regression from Boundary
We computed this as the difference in total bias between the original constrained model and a model with extended boundaries. In the model with extended boundaries, we extended the stimulus space from $[0,1]$ to a larger interval [1,2][-1,2] to remove boundary effects. The encoding function $F(p)$ was extended by continuing the resources constantly from its values at the endpoints $p=0$ and $p=1$.
Likelihood Repulsion
We computed this component as the bias arising under the same model with the extended stimulus space and the prior replaced by a uniform prior. In this model, both prior attraction and boundary regression are removed, isolating the repulsion effect.
Prior Attraction
We computed it as the difference between the mode of the posterior distribution and the expected value of the sensory likelihood function.
In Figure 33, the decomposition reveals two critical insights regarding the origin of probability distortion. First, regarding the failure of Uniform Encoding, the most distinct qualitative difference appears in the Likelihood Repulsion column. The model with uniform encoding exhibits zero likelihood repulsion. Consequently, to approximate the empirical S-shaped bias, it is forced to rely exclusively on Boundary Regression and a heavily distorted Prior Attraction, making it structurally insufficient to generate the smooth, global S-shaped distortion observed in human data. Second, the trade-off between bias and variability fit. The Gaussian prior produces a smooth attraction curve. While this fits the general shape of the bias well, the rigidity of the Gaussian form prevents it from forming the sharp peak necessary to capture the dip in response variability at $p=0.5$. The nonparametric priors exhibit fluctuations in their attraction. This distortion arises because of a sharp prior peak to capture the low variability at $p=0.5$ 1. This sharp peak exerts excessive local attraction, creating slight deviations in the bias curve to satisfy the global normalization constraint.
Figure 34 illustrates the bias decomposition for the DMR Pricing task. Similar to JRF, the UniformE model exhibits zero likelihood repulsion. It attempts to compensate for this absence by inducing a heavily distorted Prior Attraction and relying on Boundary Regression. While this compensation allows for a moderate quantitative fit in pricing (a task where we compare the expected value instead of probability), it remains mechanistically distinct from the repulsion in U-shaped models, ann we seek for more evidence in Figure 35, the bias decomposition for the DMR Choice task. As in other tasks, the Uniform model shows zero likelihood repulsion. However, unlike in the Pricing task, the Prior Attraction here remains relatively flat and fails to compensate for the missing repulsion. This flatness likely occurs because, in a binary choice task, increasing prior precision (to reduce variance) without sufficient sensory discriminability at the boundaries would only compress values and reduce the model’s ability to predict preference reversals. The model thus defaults to a smoother prior to avoid introducing excessive bias error.
Figure 36 shows how the Bayesian model captures the shift in behavior under bimodal stimulus statistics. In Prior Attraction, the GaussianP model produces a simple, smooth attraction curve. In contrast, both the FreeP and BimodalP models exhibit multi-peaked attraction patterns. These fluctuations correspond to attraction forces directed toward the two modes of the bimodal stimulus distribution.
Figure 33: JRF Task: The comparison between bias decomposition of four Bayesian model variants in Figure 3. The solid line shows the group median, and the shaded area indicates the interquartile range (IQR). Plotting details are provided in Appendix C.14.
Figure 34: DMR Pricing Task: Bias decomposition of four Bayesian model variants in Figure 5B. The solid line shows the group median, and the shaded area indicates the interquartile range (IQR). Plotting details are provided in Appendix C.14.
Figure 35: DMR Choice Task: Bias decomposition of two Bayesian model variants in Figure 5D. The solid line shows the group median, and the shaded area indicates the interquartile range (IQR). Plotting details are provided in Appendix C.14.
Figure 36: Adaptation Task: Bias decomposition of three Bayesian model variants in Figure 6D. The solid line shows the group median, and the shaded area indicates the interquartile range (IQR). Plotting details are provided in Appendix C.14.
C.14 Details for Plotting Group-level Prior and Resources
@@
C.15 Methods for Calculating Bias and Variance
@@ C.15.3 Bayesian Model
C.16 Experiments on explaining Probability Distortion in Vision-Language Models
@@ C.16.1 Task and Stimuli
C.16.2 Behavioral Findings: S-shaped Distortion
We first examined the end-to-end behavior of open-source VLMs when prompted with dot-array images. Both InternVL3-14B and Qwen-VL2.5-7B produced probability estimates showing the classic S-shaped distortion observed in humans (Figures 37, 38). Specifically, models overestimated small probabilities, underestimated large probabilities, and showed a dip in accuracy near $p=0.5$, accompanied by relatively low variability—a pattern similar to human data. These results held across both long and short prompts (Figures 39, 40).
@@ C.16.2 Behavioral Findings: S-shaped Distortion
C.16.3 Representation Analysis with A Proxy for Fisher Information
@@
We applied this method to hidden states in the decoder part of each VLM, where vision and text features interact. Both InternVL3-14B and Qwen-VL2.5-7B exhibited robust U-shaped Fisher information profiles (Figures 42, 43). The U-shape was evident from the earliest fusion layers and remained stable through the final layers, with higher discriminability near $p=0$ and $p=1$. This pattern parallels our theoretical prediction that a U-shaped allocation of resources underlies S-shaped bias.
@@ B Layer 48
C.16.4 Summary and Theoretical Implications
Together, these results show that S-shaped probability distortion also emerges in large VLMs when performing the JRF task.
This phenomenon aligns with the framework of stewart2006decision, which demonstrates that naturally occurring probability expressions in language corpora follow a distribution clustered at extremes rather than a uniform distribution. Since VLMs are trained on vast datasets reflecting these ecological statistics, they likely internalize this extreme-heavy prior. This interpretation is further supported by recent findings from zhu2025language, who explicitly showed that LLMs exhibit probability weighting patterns precisely when pretrained on data following such ecological distributions.
In our experiments, probing hidden representations reveals a U-shaped Fisher Information, consistent with the Bayesian account that nonuniform allocation of encoding resources underlies the observed S-shaped bias.